OpenStudy (anonymous):

Angle θ lies in the second quadrant and sin θ = 3/5. cos θ= a. -4/5 b. -3/5 c. 3/5 d. 4/5 Tan θ= a. -4/3 b. -3/4 c. 3/4 d. 4/3

OpenStudy (anonymous):

i think that cos= -4/5 and tan= -3/4

OpenStudy (anonymous):

You have them both right

OpenStudy (anonymous):

Ok, thank you. :)